find the cost of painting the four walls of a room 10 meters long, 5 meters broad and 6 meters high at the cost of rs 4 per square metere
Answer by TimothyLamb(4379)

---
NOTE TO STUDENT:
calculate the surface area of the four walls
---
two long walls: 10*6 sq.m
two short walls: 5*6 sq.m
---
a = 2(10*6) + 2(5*6) sq.m
a = 180 sq.m
---
cost = (area units a) * (cost per area unit r)
c = a*r
c = 180*4
c = 720 rupee
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php